# Improving Retrieval Augmented Language Model with Self-Reasoning

### v2

@@ 5.4 Main Results
In the context of long-form QA evaluations, for the metrics of EM recall, it needs to comprehend multiple documents and merge answers. The EAP and TAP are specifically designed for multi-document reading comprehension, enabling our performance to surpass other baselines. In terms of citation evaluation metrics, our self-reasoning framework can achieve better results than GPT-4 in ASQA citation recall metrics (72.3 vs. 68.5).
This is largely due to the reasoning trajectories generated in the EAP, which can enhance the recall and precision of citation evaluation, leading to more interpretable generations.
For fact verification evaluations, we observed that self-seasoning is dominantly superior to all baseline models. Our method achieves a much higher accuracy rate than the Self-RAG model (83.9 vs. 72.1). The RAP in our framework is designed to judge the relevance between the retrieved documents and the question, which leads to the notable enhancement in accuracy for this fact verification task.
To clearly demonstrate the practical applications and benefits of our self-reasoning framework, we provide a case study for a more in-depth analysis in the Appendix §$\S$§A.6, which illustrates how the framework operates in real-world scenarios.
@@
6 Analysis
6.1 Ablation Study
@@
We conduct an ablation study on two short-form QA datasets and a fact verification dataset to analyze the individual contributions of each process within our proposed self-reasoning framework. We further explore the effectiveness of the gradual learning (GL) method and the quality control (QC) of data generation (a detailed analysis described in Appendix §$\S$§A.7). The main ablation study results are shown in Table 2.
Effectiveness of Relevant-Aware Process
@@
First, we evaluate the effect of the Relevant-Aware Process (RAP). The removal of the RAP causes the overall performance to drop in two short-form QA datasets and a fact verification dataset, suggesting that preliminary consideration of the relevance between questions and retrieved documents can help improve performance. We notice that the performance declines most significantly in the FEVER dataset. Detecting irrelevant documents is critical in the fact-verification task. Our model will immediately output NotEnoughInfo if it detects that all documents are irrelevant.
Effectiveness of Evidence-Aware Process
@@
Then we evaluate the effect of the Evidence-Aware Selective Process (EAP). Removing the EAP causes the overall performance of the average accuracy to decline from 60.9 to 56.3 in three short-form QA datasets. This reduction indicates that snippets of key sentences and document citations generated through self-reasoning are instrumental in boosting accuracy.
Effectiveness of Trajectory Analysis Process
@@
Finally, we evaluate the effect of the Trajectory Analysis Process (TAP). When excluding the TAP, we can observe a performance decline on all three datasets, demonstrating that self-analysis based on two previous processes generated trajectories can also improve the performance of LLMs. Note that the analysis content generated by TAP is indispensable for the long-form QA evaluation.
6.2 Retrieval Robustness Analysis
@@
As retrievers are not perfect and past work has shown that noisy retrieval can have negative effects on the performance of LLMs Petroni et al. (2020); Li et al. (2023). In this section, we design two kinds of settings to validate the robustness of RALMs. In the first setting, we test whether the order of the retrieved documents will affect the performance of the RALMs. Specifically, after retrieving the top-$k$ documents using retrievals (such as the DPR) with a descending relevance score, we randomly shuffle the order of the retrieved documents and then input them to an LLM. In the second setting, we test how noisy documents impact the performance of LLMs. When retrieving the top-$k$ documents from the given question, we randomly replace 50% of the retrieved documents with other documents sampled from a different question in the dataset.
@@
Figure 3 shows the noise robustness experiment results on three datasets. Our self-reasoning framework consistently outperforms the Self-RAG and Vicuna models. We observe that random shuffling of retrieved documents has a minimal impact on the performance of RALMs. If the provided documents are supportive, it is trivial for a RALM to determine the correct answer. However, when presented with noisy documents, all models experience a decline in performance. The performance drop for our self-reasoning framework is relatively minimal, which demonstrates that our proposed method is robust even when dealing with noisy documents.
@@
6.3 Citation analysis
As the automatic evaluation by the NLI model cannot detect partially supported citations, we discuss the analysis of citations with human evaluation in this section. Similarly to Liu et al. (2023), we conduct a human evaluation on two dimensions: 1) citation recall: annotators are given a statement and all documents that the statement refers to and are asked to judge whether the documents fully support the given statement; 2) citation precision: given a statement and one of its citations, annotators are asked to validate whether the citation fully supports, partially supports or does not support the statement. Each citation gets a precision score of 1 if the output sentence has a citation recall of 1 and this citation at least partially supports it. Details of human annotation can be found in Appendix §$\S$§A.8.
As shown in Figure 4, we observe that the relative rankings by human evaluation align well with those from the automatic evaluation, and the human evaluation often yields a closely higher score when compared with the automatic evaluation.
7 Conclusion
@@
8 Limitations
In this work, we mainly focus on improving the performance of RALMs with a self-reasoning framework on the task of open domain question answering and fact verification. Although we believe our framework can involve the distribution of real-world user questions, as we have evaluated in four public datasets, we acknowledge that we have not explored more challenging scenarios, such as multi-hop reasoning, code generation, and arithmetic reasoning. In future work, more challenging reasoning tasks, such as arithmetic reasoning, should be explored for the self-reasoning framework. We believe that our framework can effectively mitigate factual hallucinations in LLMs and improve the robustness of RALMs. However, there is still a risk that our method might generate hallucinations.
@@ Appendix A Appendix
Self-RAG Asai et al. (2024) N N N No Need 145K (Generator)/ 46K (Critic)
ReAct Yao et al. (2023) N N Y No Need No Need
RECOMP Xu et al. (2023a) N Y Y 152K No Need
@@
A.1 More Related Work of LMs for Reasoning
One of the most well-known methods of using LLMs for reasoning is the Chain-of-Thought (CoT) Wei et al. (2022), which demonstrates the capability of LLMs to create their thinking process for problem-solving. Zhou et al. (2022) proposes a least-to-most prompting for solving complex tasks. Wang et al. (2022) introduces a method to reason with self-consistency. Press et al. (2023) proposes a method to further improve the chain of thought by reasoning explicitly instead of implicitly.
Recent works have extended beyond the internal reasoning ability of LLMs to include interactions with external tools (e.g., search engines or retrievers) for solving complex tasks. The ReAct Yao et al. (2023) presents an iterative paradigm to combine reasoning and acting with LLMs for tackling language reasoning and decision-making tasks. Xu et al. (2023b) introduces a framework to enable information retrieval and LLMs to interact with each other effectively with chain-of-query decomposition. Pan et al. (2024) proposes a novel framework named Chain-of-Action (CoA), which integrates a reasoning retrieval method to decompose complex questions into chains of configurable actions.
@@ A.3 Datasets Description
NaturalQuestion (NQ) Kwiatkowski et al. (2019) contains real user questions issued to the Google search and answers found from Wikipedia by the annotators. NQ is created to train and evaluate automated question answering systems.
PopQA Mallen et al. (2023) is a large-scale open-domain question answering dataset, consisting of entity-centric QA pairs. Each question is made by converting a knowledge triplet retrieved from Wikidata using a template. In this work, we use PopQA to evaluate performance in long-tail settings.
ASQA Stelmakh et al. (2022) is a long-form factoid dataset, and most questions can be answered by Wikipedia. Each question originates from AmbigQA Min et al. (2020) and represents an ambiguous query that requires multiple short answers to cover various aspects. The dataset provides a long-form answer that contains all short answers.
FEVER Thorne et al. (2018) is a fact verification dataset that contains claims generated by rewriting sentences extracted from Wikipedia and subsequently verified without knowledge of the sentence from which they were derived. The claims are classified as Supported, Refuted, or NotEnoughInfo.
A.4 Experiment settings
@@
During gradual learning, we fine-tune the LLaMA-2 Touvron et al. (2023) model with our self-reasoning framework for 3 epochs with a batch size set to 32, leveraging the DeepSpeed library Rasley et al. (2020) and the ZeRO optimizer, and we use parameter partitioning ZeRO stage 3 with float16 precision. The learning rate $r_{a}$ for the first stage is set to 5e-5, the learning rate $r_{b}$ for the second stage is set to 3e-5, and the learning rate $r_{c}$ for the final stage is set to 1e-5. Our self-reasoning 13B model is trained on the NVIDIA Tesla 8 $\times$ V100 32GB GPU for 4 hours, while the 7B model is trained for 2 hours.
@@
We use the vLLM  [^2] Codes are available at https://github.com/vllm-project/vllm framework Kwon et al. (2023) to accelerate the inference speed during inference. The codes follow the Apache-2.0 license agreement. We use greedy decoding in all experiments to ensure deterministic generations. We test the temperature within a range of $\{0.2,0.4,0.6,0.8,1.0\}$, finally we set the temperature to 0.2, as we observed lower temperature results in better performance in the open-domain question answering task. The maximum generation length is set to 2048 for our model. All baseline models are tested with zero-shot settings for short-form QA datasets, and with one-shot settings for the long-form QA and fact verification datasets.
@@
A.5 Categorical Comparisons
We differentiate our method from existing strong baseline models by categorizing and comparing it across six dimensions, as presented in Table 3. As illustrated in the table, our method can stand out in several key aspects.
@@ A.5 Categorical Comparisons
A.6 Case Study
@@
A.7 More Analysis on Ablation Study
Effectiveness of Gradual Learning
@@ A.7 More Analysis on Ablation Study
Effectiveness of Quality Control
@@
A.8 Human Evaluation
We randomly sample 100 examples from the ASQA dataset and annotate the outputs of selected models. Each sample is then assigned to two people for annotation. Each annotator is required to verify the citation recall and citation precision acorrding to our provided scheme. The annotation scheme is inspired by Gao et al. (2023b) as follows:

### v3

@@ Introduction
In the context of long-form QA evaluations, for the metrics of EM recall, it needs to comprehend multiple documents and merge answers. The EAP and TAP are specifically designed for multi-document reading comprehension, enabling our performance to surpass other baselines. In terms of citation evaluation metrics, the self-reasoning RAG can achieve better results than GPT-4 in ASQA citation recall metrics (72.3 vs. 68.5). This is largely due to the reasoning trajectories generated in the EAP, which can enhance the recall and precision of citation evaluation, leading to more interpretable and traceable generations.
For fact verification evaluations, we observed that self-seasoning is dominantly superior to all baseline models. Our method achieves a much higher accuracy rate than the Self-RAG model (83.9 vs. 72.1). The RAP in our framework is designed to judge the relevance between the retrieved documents and the question, which leads to a notable enhancement in accuracy for this fact verification task.
To clearly demonstrate the practical applications and benefits of our self-reasoning framework, we provide a case study for a more in-depth analysis in Appendix, which illustrates how our framework operates in real-world scenarios.
@@
Models NQ PopQA FEVER
(acc) (acc) (acc)
LLaMA2 32.7 53.5 53.4
+ trajectory 38.3 54.2 79.2
[…]
+ trajectory 38.5 56.4 79.6
Table 3: The analysis on the effectiveness of self-reasoning trajectories with 13B parameter size models. In the table, the +trajectory indicates the result of the baseline model is enhanced with self-generated trajectories by our framework.
Analysis
Ablation Study
@@
We conduct an ablation study on two short-form QA datasets and a fact verification dataset to analyze the individual contributions of each process within our proposed self-reasoning framework. We further explore the effectiveness of the gradual learning (GL) method and the quality control (QC) of data generation (a detailed analysis described in the Appendix). The main ablation study results are shown in Table 2 and Table 3.
Effectiveness of RAP.
@@
First, we evaluate the effect of the RAP. The removal of the RAP causes the overall performance to drop in two short-form QA datasets and a fact verification dataset, suggesting that preliminary consideration of the relevance between questions and retrieved documents can help improve performance. We notice that the performance declines most significantly in the FEVER dataset. Detecting irrelevant documents is critical in the fact-verification task. Our model will immediately output NotEnoughInfo if it detects that all documents are irrelevant.
Effectiveness of EAP.
@@
Then we evaluate the effect of the EAP. Removing the EAP causes the overall performance of the average accuracy to decline from 60.9 to 56.3 in three short-form QA datasets, which indicates that snippets of key sentences and document citations generated through self-reasoning are instrumental in boosting accuracy.
Effectiveness of TAP.
@@
Finally, we evaluate the effect of the TAP. When excluding the TAP, we can observe a performance decline on all three datasets, demonstrating that self-analysis based on two previous processes generated trajectories can also improve the performance of LLMs. Note that the analysis content generated by TAP is indispensable for the long-form QA evaluation.
Effectiveness of Self-Reasoning Trajectory.
To verify whether the trajectories generated by the self-reasoning framework are truly effective, we put the trajectories generated by our self-reasoning framework into the original baseline models as input prompts, and then use the baseline models to regenerate the answers. We observe that incorporating self-generated trajectories can significantly enhance performance in short QA tasks and fact verification tasks.
Retrieval Robustness Analysis
@@
Retrievers are not perfect and past work has shown that noisy retrieval can have negative effects on the performance of LLMs (Petroni et al. 2020; Li et al. 2023). In this section, we design two kinds of settings to validate the robustness of RALMs. In the first setting, we test whether the order of the retrieved documents will affect the performance of the RALMs. Specifically, after retrieving the top-$k$ documents using retrievals with a descending relevance score, we randomly shuffle the order of the retrieved documents and then input them to an LLM. In the second setting, we test how noisy documents impact the performance of LLMs. When retrieving the top-$k$ documents from the given question, we randomly replace 50% of the retrieved documents with other documents sampled from a different question in the dataset.
@@
Figure 3 shows the noise robustness experiment results on three datasets. Our self-reasoning framework consistently outperforms the Self-RAG and Vicuna models. We observe that random shuffling of retrieved documents has a minimal impact on the performance of RALMs. If the provided documents are supportive, it is trivial for a RALM to determine the correct answer. However, when presented with noisy documents, all models experience a decline in performance. The performance drop in our self-reasoning framework is relatively minimal, demonstrating the robustness of our method even when handling noisy documents.
@@
Citation Analysis
As the automatic evaluation by the NLI model cannot detect partially supported citations, we discuss the analysis of citations with human evaluation in this section. Similarly to Liu, Zhang, and Liang (2023), we conduct a human evaluation on two dimensions: 1) citation recall: annotators are given a statement and all documents that the statement refers to and are asked to judge whether the documents fully support the given statement; 2) citation precision: given a statement and one of its citations, annotators are asked to validate whether the citation fully supports, partially supports or does not support the statement. As shown in Figure 4, the relative rankings by human evaluation align well with those from the automatic evaluation, and the human evaluation often yields a closely higher score when compared with the automatic evaluation. Details of human annotation can be found in the Appendix.
Latency Analysis
We also compared the inference latency of self-reasoning RAG with that of Self-RAG and GPT-4. The results show that our method maintains comparable latency to Self-RAG while delivering substantial performance gains. Detailed results are available in the Appendix.
Conclusion
@@ Appendix A Appendix
Self-RAG (Asai et al. 2024) N N N No Need 145K (Generator)/ 46K (Critic)
ReAct (Yao et al. 2023) N N Y No Need No Need
RECOMP (Xu et al. 2024) N Y Y 152K No Need
@@
More Related Work of LMs for Reasoning
One of the most well-known methods of using LLMs for reasoning is the Chain-of-Thought (CoT) (Wei et al. 2022), which demonstrates the capability of LLMs to create their thinking process for problem-solving. Zhou et al. (2022) proposes a least-to-most prompting for solving complex tasks. Wang et al. (2022) introduces a method to reason with self-consistency. Press et al. (2023) proposes a method to further improve the chain of thought by reasoning explicitly instead of implicitly.
Recent works have extended beyond the internal reasoning ability of LLMs to include interactions with external tools (e.g., search engines or retrievers) for solving complex tasks. The ReAct (Yao et al. 2023) presents an iterative paradigm to combine reasoning and acting with LLMs for tackling language reasoning and decision-making tasks. Xu et al. (2023) introduces a framework to enable information retrieval and LLMs to interact with each other effectively with chain-of-query decomposition. Pan et al. (2024) proposes a novel framework named Chain-of-Action (CoA), which integrates a reasoning retrieval method to decompose complex questions into chains of configurable actions.
@@
NaturalQuestion (NQ) (Kwiatkowski et al. 2019) contains real user questions issued to the Google search and answers found from Wikipedia by the annotators. NQ is created to train and evaluate automated question answering systems.
PopQA (Mallen et al. 2023) is a large-scale open-domain question answering dataset, consisting of entity-centric QA pairs. Each question is made by converting a knowledge triplet retrieved from Wikidata using a template. In this work, we use PopQA to evaluate performance in long-tail settings.
ASQA (Stelmakh et al. 2022) is a long-form factoid dataset, and most questions can be answered by Wikipedia. Each question originates from AmbigQA (Min et al. 2020) and represents an ambiguous query that requires multiple short answers to cover various aspects. The dataset provides a long-form answer that contains all short answers.
FEVER (Thorne et al. 2018) is a fact verification dataset that contains claims generated by rewriting sentences extracted from Wikipedia and subsequently verified without knowledge of the sentence from which they were derived. The claims are classified as Supported, Refuted, or NotEnoughInfo.
Experiment settings
@@
During gradual learning, we fine-tune the LLaMA-2 (Touvron et al. 2023) model with our self-reasoning framework for 3 epochs with a batch size set to 32, leveraging the DeepSpeed library (Rasley et al. 2020) and the ZeRO optimizer, and we use parameter partitioning ZeRO stage 3 with float16 precision. The learning rate $r_{a}$ for the first stage is set to 5e-5, the learning rate $r_{b}$ for the second stage is set to 3e-5, and the learning rate $r_{c}$ for the final stage is set to 1e-5. Our self-reasoning 13B model is trained on the NVIDIA Tesla 8 $\times$ V100 32GB GPU for 4 hours, while the 7B model is trained for 2 hours.
@@
We use the vLLM  [^2] Codes are available at https://github.com/vllm-project/vllm framework (Kwon et al. 2023) to accelerate the inference speed during inference. The codes follow the Apache-2.0 license agreement. We use greedy decoding in all experiments to ensure deterministic generations. We test the temperature within a range of $\{0.2,0.4,0.6,0.8,1.0\}$, finally we set the temperature to 0.2, as we observed lower temperature results in better performance in the open-domain question answering task. The maximum generation length is set to 2048 for our model. All baseline models are tested with zero-shot settings for short-form QA datasets, and with one-shot settings for the long-form QA and fact verification datasets.
@@
Categorical Comparisons
We differentiate our method from existing strong baseline models by categorizing and comparing it across five dimensions, as presented in Table 4. As illustrated in the table, our method can stand out in several key aspects.
@@
Algorithm 1 Data Quality Control
1:the origin self-reasoning dataset $\mathcal{D}_{0}$ generated by GPT-4
2:filtered high quality self-reasoning dataset $\mathcal{D}_{sr}$
3:Initialize the evaluation metrics program $P$
4:Initialize the citation score tools $T$ (Gao et al. 2023b)
5:for $i=1$ to $N$ do
6:     Evaluate whether the answer of sample $d_{i}$ is correct using program $P$
7:     if True and $d_{i}\in$ Long-form QA dataset  then
8:         Compute citation recall score $s_{r}$ for sample $d_{i}$
9:         Compute citation precision score $s_{p}$ for sample $d_{i}$
10:         if  $s_{r}\geq\delta_{r}$ and $s_{p}\geq\delta_{p}$  then
11:              Add the training sample to $\mathcal{D}_{l}$
12:         end if
13:     end if
14:     if True and $d_{i}\notin$ Long-form QA dataset  then
15:         Add the training sample to $\mathcal{D}_{s}$
16:     end if
17:end for
18:$\mathcal{D}_{sr}=\mathcal{D}_{l}\cup\mathcal{D}_{s}$
19:return $\mathcal{D}_{sr}$
Case Study
@@
More Analysis on Ablation Study
Effectiveness of Gradual Learning.
@@
Effectiveness of Quality Control.
@@
Human Evaluation
We randomly sample 100 examples from the ASQA dataset and annotate the outputs of selected models. Each sample is then assigned to two people for annotation. Each annotator is required to verify the citation recall and citation precision acorrding to our provided scheme. The annotation scheme is inspired by (Gao et al. 2023b) as follows:
@@
The Pseudo-code of Data Quality Control
The details and pseudo-code of data quality control are illustrated in Algorithm 1. In the table, $\mathcal{D}_{0}$ is the origin self-reasoning dataset generated by GPT-4. $P$ is the program to evaluate the answer according to the metrics. $T$ is the off-the-shelf tool (Gao et al. 2023b) to calculate the citation scores. $N$ is the number of total samples from the origin dataset, and $d_{i}$ is the $i$-th training sample. $s_{r}$ and $s_{c}$ are citation recall and precision scores for each sample respectively. $\delta_{r}$ and $\delta_{p}$ are thresholds of citation recall and precision respectively. $\mathcal{D}_{l}$ is filtered trajectory data for long-form QA, and $\mathcal{D}_{s}$ is filtered dataset for short-form QA and fact verification dataset. $\mathcal{D}_{sr}$ is the final high quality self-reasoning training dataset.
@@
Latency Analysis
We conduct extra experiments to measure inference latency across our approach, Self-RAG, and GPT-4. The results demonstrated that our method achieved better performance while maintaining latency comparable to Self-RAG. The results of average inference latency per question for the NQ and ASQA datasets are shown in Table 5.
Models NQ ASQA
LLaMA2 0.19 1.92
Self-RAG 2.08 2.38
Self-Reasoning 2.19 2.32
[…]
Table 5: The latency experiment results on NQ and ASQA datasets. The average inference latency per question is evaluated using 7B-parameter models.